# SAM 2++: Tracking Anything at Any Granularity

### v3

@@ 10.1 Performance Comparison
Unicorn [Unicorn] 69.2 65.2 73.2 83.0 86.4 82.2
UNINEXT-H [UNINEXT] 81.8 77.7 85.8 85.4 89.0 86.4
MITS [MITS] 84.9 82.0 87.7 83.4 88.9 84.6
OmniTracker-L [OmniTracker] 71.0 66.8 75.2 83.4 86.7 82.3
@@
Comparison with Unified Models. To demonstrate the superior performance of our unified tracking model, we conducted a comparative evaluation against other unified models. As shown in Table 13, our model achieves significantly better results on two classical benchmark datasets, highlighting its remarkable effectiveness and robustness.
Table 14: Comparison with TAG Model.
Methods VOST LaTOT TOTB CroHD
TAG [harley2024tag] 31.3 35.3 74.4 57.1
[…]
Comparison with TAG model. We found that there exists a work, TAG [harley2024tag], that shares the same objective as ours, which is to achieve tracking of mask, box, and point with a unified model. To highlight the contribution of our work, we provide a detailed comparison between the two approaches. Firstly, TAG is an offline tracking model that processes multiple frames as a clip simultaneously, which not only differs from the current mainstream online frame-by-frame tracking pipeline but also leads to information leakage from future frames and is only applicable to pre-recorded videos. In contrast, our model ensures that the current frame only receives information from the past frames, making it suitable for video streams. Secondly, in the way of the prompt construction, TAG simply converts point coordinates into a {0, 1} mask in the point task, providing limited target information. Our method combines point coordinates with a (0, 1) Gaussian mask: the former provides precise locations, while the latter highlights the target point in mask form, maintaining consistency with output from MaskDecoder and input for MemoryEncoder, thereby enhancing expressiveness. For the box task, TAG converts the box into a square mask, which causes confusion between the target region and the background, affecting the accuracy of the target information. Third, TAG is trained only on public datasets, which limits the scale of the training dataset. In contrast, we construct a Data Engine that enables both model training and dataset annotation expansion, ultimately resulting in a large-scale dataset with three types of granularity annotations and a well-trained model. Most important of all, as an offline approach, the TAG model primarily focuses on how to jointly encode targets of varying granularity. When processing the next clip, the prompt remains in an original, unmodeled form, lacking rich target representation (e.g., mask, point, or box). Meanwhile, due to the lack of judgment of predictions, the next clip must adopt the prediction of the last frame in the previous clip as a prompt, even if it may be unreliable, which leads to error accumulation and makes it difficult to handle common challenges such as temporary target disappearance. In contrast, our method follows the online setting. The core challenges lie not only in multi-granularity prompts encoding, but also in how to transform predictions into memory representations to guide subsequent frames. Compared to original prompts, memory offers richer target features that improve the stability and accuracy. Leveraging both selective capability and memory diversity, the mechanism compensates for potential errors in individual predictions, effectively improving stability. To make this key component compatible with multi-granularity tracking, we introduced a Task-Adaptive Memory mechanism, one of the major contributions, to unify the predictions of varying granularity into memory representations. In summary, the distinction is not merely at the task level, but it directly impacts motivation and core innovations. Additionally, we present a performance comparison on the three tasks. As shown in the Table 14, our model significantly outperforms the TAG model on all three tasks, demonstrating the superior performance of our model.
Table 15: Analysis of training data and task mixtures on three tracking tasks.
@@
Qualitative Results. We first demonstrate the multi-granularity tracking capabilities of SAM 2++ across multiple benchmarks, as illustrated in Fig. 8. We further compare our method qualitatively with various SOTA models in three tasks. As shown in Fig.9, our model outperforms other models across all three tasks. This demonstrates that our model effectively handles various target state granularities while also exhibiting strong robustness and generalization to diverse scenarios and challenges.
10.2 Model and Data Ablation
Study on model setting of point tracking task. We compare the performance of point tracking under different model settings in Table 16. First, performance declines when the Gaussian mask prompt for the point tracking task is removed, indicating that incorporating the Gaussian mask effectively assists the mask output of the Decoder, and demonstrating the effectiveness of our proposed task-specific prompt. Second, we compare two approaches for obtaining point coordinates: applying argmax to the mask output v.s. adding an MLP to predict the coordinates directly. The results show that the argmax operation yields better performance, suggesting that argmax is an effective method for point prediction, supervises the mask output as a memory source, and better represents the target state at the point granularity. In contrast, the additional MLP requires adaptation to the original model and struggles to supervise the mask output effectively.
Table 16: Analysis of the model setting on point tracking task.
@@ 10.2 Model and Data Ablation
[…]
Study on task mixture and training data. We compare the performance under different training settings for three tracking tasks in Table 15. For evaluating original SAM 2 on the single object task, we take the ground-truth bounding box from the first frame as a box prompt to predict the target mask, then predict the mask frame by frame, and finally extract the outer bounding box from each mask as the final box prediction. After training on the public dataset and Phase ① of our Tracking-Any-Granularity dataset, the performance of our SAM 2++ model improves across all three tasks, demonstrating the advantages of our model design. More importantly, when we further incorporate two additional tasks during training, the model’s performance on both tasks surpasses that of training on a single task alone. This illustrates two core motivations behind our proposed model: (1) Although the granularity of the target states in the three tasks differs, they all can adopt the ”matched memory” tracking paradigm. Thus, training on various tasks enhances the matching ability, which in turn improves the performance of all tracking tasks. (2) As a generalized model supporting multiple tasks, SAM 2++ can be trained on large-scale datasets for multiple tasks, rather than being restricted to individual tasks. Finally, under the task-mixed training setting, incorporating our proposed dataset further improves the model performance on both tasks. This improvement demonstrates that the diverse and comprehensive annotations included in our dataset provide valuable supervision signals for the model, enabling it to learn more robust and generalizable representations.
@@ 11 Limitations, Impacts, and Future
However, the model still has some limitations. First, the current version does not yet support language- and audio-based references. Addressing this limitation requires integrating corresponding feature extractors into the Prompt Encoder to accommodate more types of reference states, as well as introducing relevant datasets for training. Second, in our task-specific memory, some parameters of the memory-related modules are decoupled for different tasks. Although this mechanism only adds a minimal number of parameters, these parameters are supervised by a single task and cannot benefit from multi-task learning as the majority of shared parameters do. To address the issues caused by decoupled parameters, one approach is to employ an adapter that unifies memory across different granularities, another is to fuse the decoupled parameters and dynamically adjust their scaling according to the specific task. Additionally, SAM 2++ still faces challenges in accurately tracking objects under severe occlusion, fast motion, and the presence of similar distractors. To further enhance model performance in these difficult scenarios, introducing motion modeling mechanisms and specialized memory designs could be effective solutions.

### v4

@@ 6.1 Comparison to state-of-the-art on three tasks
Unicorn [75] 69.2 65.2 73.2 83.0 86.4 82.2
UNINEXT-H [76] 81.8 77.7 85.8 85.4 89.0 86.4
MITS [74] 84.9 82.0 87.7 83.4 88.9 84.6
OmniTracker-L [62] 71.0 66.8 75.2 83.4 86.7 82.3
Ours 89.1 86.3 91.9 86.0 90.1 87.3
Comparison with Unified Models. To demonstrate the superior performance of our unified tracking model, we conducted a comparative evaluation against other unified models. As shown in Table X, our model achieves significantly better results on two classical benchmark datasets, highlighting its remarkable effectiveness and robustness.
Qualitative Results. We first compare our method with representative SOTA models across three tracking tasks. As shown in Fig. 6, SAM 2++ produces more stable and accurate results across mask, box, and point tracking, demonstrating robustness and generalization to diverse scenarios. We further visualize the multi-granularity tracking capability of SAM 2++ across multiple benchmarks in Fig. 7, where the same unified model handles different target state granularities with correct predictions.
TABLE XI: Analysis of Mixed training strategy and Data Engine.
‘Shared’ and ‘Decoupled’ denote that the module shares parameters or decouples parameters between different tasks, respectively.
@@
$\times$ Shared Shared 69.3 84.7 28.1 26.1 64.1 30.3
$\times$ Shared Decoupled 73.6 86.4 64.5 57.3 63.0 28.7
+① Freezed Decoupled 74.2 87.1 67.8 59.2 68.8 29.9
+① Decoupled Decoupled 75.4 86.8 66.9 58.5 72.1 32.7
+① Shared Decoupled 76.4 87.1 68.9 58.2 71.9 33.1
+①,②,③ Shared Decoupled 77.8 87.4 70.6 59.2 72.8 35.2
TABLE XII: Analysis of training data and task mixtures on three tracking tasks. The table compares single-task and mixed-task training settings, showing how joint training and additional TAG data affect performance across mask, box, and point tracking.
@@
#3 74.7 70.6 78.8 86.7 95.4 88.6 70.9 78.3 76.7 71.9 81.4
6.2 Exploration Studies
Study on Mixed training strategy. To verify the effectiveness of task-adaptive memory during multi-task joint training, we compare the results of single-task training with different parameter settings during multi-task mixing training. As shown in Table XI, when a single set of parameters is naively shared for multi-task joint training, the differences between tasks lead to a performance drop across all tasks. This indicates that the encoding and retrieval components of the memory module need to be decoupled for different tasks. In addition, when the image encoder is either frozen or similarly decoupled, the performance is inferior to a shared encoder. This suggests that the image encoder benefits from exposure to more data and is not adversely affected by task differences.
Study on data engine. To validate the effectiveness of our data engine, we evaluated the performance when trained on different phases of our TAG dataset, as shown in Table XI. The results demonstrate that our proposed dataset enhances the performance on other datasets, indicating high diversity and generalizability. After training with more data from phases ② and ③, the performance is further improved across all three tasks, demonstrating the effectiveness of the supplementary data provided by our data engine.
Study on task mixture and training data. We compare the performance under different training settings for three tracking tasks in Table XII. For evaluating original SAM 2 on the single object task, we take the ground-truth bounding box from the first frame as a box prompt to predict the target mask, then predict the mask frame by frame, and finally extract the outer bounding box from each mask as the final box prediction. After training on the public dataset and Phase ① of our Tracking-Any-Granularity dataset, the performance of our SAM 2++ model improves across all three tasks, demonstrating the advantages of our model design. More importantly, when we further incorporate two additional tasks during training, the model’s performance on both tasks surpasses that of training on a single task alone. This illustrates two core motivations behind our proposed model: (1) Although the granularity of the target states in the three tasks differs, they all can adopt the ”matched memory” tracking paradigm. Thus, training on various tasks enhances the matching ability, which in turn improves the performance of all tracking tasks. (2) As a generalized model supporting multiple tasks, SAM 2++ can be trained on large-scale datasets for multiple tasks, rather than being restricted to individual tasks. Finally, under the task-mixed training setting, incorporating our proposed dataset further improves the model performance on both tasks. This improvement demonstrates that the diverse and comprehensive annotations included in our dataset provide valuable supervision signals for the model, enabling it to learn more robust and generalizable representations.
To make a fairer comparison, we added the results for the mask and bounding box tasks using separate public datasets for each task to Table XIII (point tracking is excluded because keypoint data is insufficient for individual training). Our model still outperforms the dedicated baselines, demonstrating that our model’s performance advantage lies not only in its ability to utilize multi-task datasets but also in the superiority of its architectural design.
TABLE XIII: Single-task comparison using public training datasets. We compare SAM 2++ trained separately for mask and box tracking with dedicated task-specific baselines under comparable public-data settings.
@@ 6.2 Exploration Studies
[…]
Ours (mask training only) 81.6 78.2 85.1
Model VastTrack
[…]
Ours (box training only) 50.1 59.6 53.2
TABLE XIV: Analysis of the model setting on point tracking task.
@@
#3 Coord. & Gauss. Mask argmaxargmax(mask) 66.2
TABLE XV: Hyperparameters and details of SAM 2++ training in three tasks.
Modules GFlops Main Param Lora Param
Image Encoder 264.4 69.1 22.25
Memory Encoder 5.0 1.4 $\times$ 3 -
Mask Decoder 53.4 9.9 -
Memory Attention 27.4 5.9 4.3 $\times$ 3
Total 350.2 89.1 35.1
Figure 8: Visualization of memory design at different components and granularities. In the visualization of each component, the left side represents our task-adaptive memory mechanism, and the right side represents full parameter sharing.
Study on model setting of point tracking task. We compare the performance of point tracking under different model settings in Table XIV. First, performance declines when the Gaussian mask prompt for the point tracking task is removed, indicating that incorporating the Gaussian mask effectively assists the mask output of the Decoder, and demonstrating the effectiveness of our proposed task-specific prompt. Second, we compare two approaches for obtaining point coordinates: applying argmax to the mask output v.s. adding an MLP to predict the coordinates directly. The results show that the argmax operation yields better performance, suggesting that argmax is an effective method for point prediction, supervises the mask output as a memory source, and better represents the target state at the point granularity. In contrast, the additional MLP requires adaptation to the original model and struggles to supervise the mask output effectively.
6.3 Efficiency Analysis
To present a comprehensive view of the model’s computational complexity and parameter overhead, we provide a detailed breakdown of the computational cost for each module in Table XV, including GFLOPs, the number of parameters, and the LoRA parameters introduced during training. We would like to kindly note that the LoRA parameters exist only during training and are merged into the main model weights at inference time. Therefore, they do not introduce additional parameters or computational overhead during inference. In addition, although the model contains multiple Memory Encoders designed for different granularities, only the branch corresponding to the current granularity is activated at inference, while the others remain inactive (excluded from computation), ensuring that no extra inference cost is incurred. In summary, the inherent inconsistency of memory is a problem that cannot be solved by a fully parameter-sharing design, while our task-adaptive memory mechanism has done its best to reduce decoupling and overhead to achieve unification.
6.4 Visualization
@@ 7 Limitations, Impacts, and Future
However, the model still has some limitations. First, the current version does not yet support language- and audio-based references. Addressing this limitation requires integrating corresponding feature extractors into the Prompt Encoder to accommodate more types of reference states, as well as introducing relevant datasets for training. Second, in our task-specific memory, some parameters of the memory-related modules are decoupled for different tasks. Although this mechanism only adds a minimal number of parameters, these parameters are supervised by a single task and cannot benefit from multi-task learning as the majority of shared parameters do. To address the issues caused by decoupled parameters, one approach is to employ an adapter that unifies memory across different granularities, another is to fuse the decoupled parameters and dynamically adjust their scaling according to the specific task. Additionally, SAM 2++ still faces challenges in accurately tracking objects under severe occlusion, fast motion, and the presence of similar distractors. To further enhance model performance in these difficult scenarios, introducing motion modeling and specialized memory designs could be effective solutions.
8 Conclusion
We present SAM 2++, a unified framework for tracking targets at any granularity. Instead of treating mask, box, and point tracking as separate problems, SAM 2++ unifies them through task-specific prompts, a Unified Decoder, and a task-adaptive memory mechanism that represents target states across granularities. To provide unified training and testing data, we also introduce the Tracking-Any-Granularity dataset, built through a data engine that combines manual annotation with model-assisted completion, providing aligned three annotations in real-world videos. Experiments across video object segmentation, single object tracking, and online point tracking show that SAM 2++ achieves strong performance at all three granularities. We hope SAM 2++ can serve as a useful baseline for general tracking and inspire future research built around more flexible target representations and unified models.
9 Acknowledgements
This work is supported by the National Key R&D Program of China (No. 2022ZD0160900), the Basic Research Program of Jiangsu (No. BK20250009), the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (No. JYB2025XDXM118), and the Collaborative Innovation Center of Novel Software Technology and Industrialization.